The healthy spring water company sells 10-gallon bottles of water for water fountains. It sells 2000 bottles a day at a price of $20. The company’s daily revenues and costs are:

            Sales revenues             $40,000

            Variable costs             $16,000

            Fixed costs                  $20,000

Although the company grew rapidly in the past decade, sales have been rather stagnant in the last year. The problem is that the market for spring water has grown large enough that grocery stores have begun to carry it at prices below those of healthy spring.

The company is considering raising the price by 15% and repositioning its water as a premium brand. Healthy Spring believes that it will be able to keep most of its loyal customers with a price increase. How many units loss can Healthy Spring tolerate before the price increase would become unprofitable?
Answer #1

A B C D E F G H I Sales VC Cont Fixed cost Profit 40000 16000 24000 20000 4000 Revised price Revised units Units loss 23 1739

A B C E F Sales VC Cont Fixed cost Profit 40000 16000 24000 20000 =E7-E8 Revised price Revised units Units loss =20*115% =E5/